Terry’s utility function over leisure (L) and other goods (Y ) is U(L, Y ) = Y + LY. The associated marginal utilities are MUY = 1 + L and MUL = Y. He purchases other goods at a price of $1, out of the income he earns from working. Show that, no matter what Terry’s wage rate, the optimal number of hours of leisure that he consumes is always the same.

(a)   What is the number of hours he would like to have for leisure?

(b)  Determine the MRS of leisure for labour

Draw a leisure-influenced labor curve


1
Expert's answer
2021-05-16T17:49:57-0400

Terry’s utility function over leisure (L) and other goods (Y ) is "U(L, Y ) = Y + LY."

The associated marginal utilities are "MU_Y = 1+L" and"MU_L = Y"

He purchases other goods at a price of $1, out of the income he earns from working. 

(a)

Here,

"MRS=\frac{MU_Y}{MU_L}"

And,

"=\frac{1+L}{Y}"

"\frac{P_Y}{P_L}=\frac{1}{W}"

So at optimality

"MRS=\frac{P_Y}{P_L}"

"\frac{Y}{1+L}=W"

"Y=W(1+L)"

now,

"WL+(1+L)=24W"

"2WL=23W"

"L=\frac{23W}{2W}"

"=11.5"

And

"Y=W(12.5)"


(b)

The MRS is

"MPS=\frac{1+L}{Y}"

"=\frac{1+11.5}{12.5W}"


"=\frac{1}{W}"

The MRS is "\frac{1}{W}"